OpenStudy (anonymous):

how do you find the answer to this? Complete the square for the equation: x2 + 2x - 3 = 0

OpenStudy (anonymous):



OpenStudy (johnweldon1993):

well yeah its EASY to factor...but to complete the square....first...HALVE the x value..what is that?

OpenStudy (johnweldon1993):

we have x² + 2x - 3 = 0 so if we HALVE the x value (not the x² value) what do we have?

OpenStudy (anonymous):

Oh ive got it now thanks!

OpenStudy (anonymous):

Oh ive got it now thanks!

OpenStudy (johnweldon1993):

okay id you just wanted it factored....then yeah that's it....I'll write down the process of completing the square anyways.... we have x² + 2x - 3 = 0 add 3 to both sides x² + 2x = 3 Halve the x value...2/2 = 1...then SQUARE THIS number....1² = 1 put this value where you see I've left the space....and also add it to the other side of the equation so x² + 2x +1 = 4 make the left hand side a square form (x+1)² = 4 take the square root of both sides x+1 = √4 x + 1 = ±2 now just solve for x....x = ±2 - 1 x = 2-1 and -2-1 x = 1 and -3 That is the process...when factoring gets harder...this makes it easier

OpenStudy (johnweldon1993):

because as you can see...it leads to the same answer....(x+1) and (x-3)

OpenStudy (anonymous):

yes i understand its just i figured it out on my paper before you asked me your question thanks anyway you were very helpful!

OpenStudy (johnweldon1993):

ahh okay....well then good job :)

OpenStudy (anonymous):

thanks(:

OpenStudy (anonymous):

i guess im really dumb at math

OpenStudy (johnweldon1993):

why do you say that? @Luis_Rivera you came to the correct answer! just you did it a different way then the question specified...still correct though